# Unraveling overoptimism and publication bias in ML-driven science

### v2

@@ Abstract
Machine Learning (ML) is increasingly used across many disciplines with impressive reported results across many domain areas. However, recent studies suggest that the published performance of ML models are often overoptimistic. Validity concerns are underscored by findings of an inverse relationship between sample size and reported accuracy in published ML models, contrasting with the theory of learning curves where accuracy should improve or remain stable with increasing sample size. This paper investigates factors contributing to overoptimistic accuracy reports in ML-driven science, focusing on data leakage and publication bias. We introduce a novel stochastic model for observed accuracy, integrating parametric learning curves and the aforementioned biases. We then construct an estimator that corrects for these biases in observed data. Theoretical and empirical results show that our framework can estimate the underlying learning curve, providing realistic performance assessments from published results. Applying the model to meta-analyses in ML-driven science, including neuroimaging-based and speech-based classifications of neurological conditions, we find prevalent overoptimism and estimate the inherent limits of ML-based prediction in each domain.
@@ 1 Introduction
The authors in [8] postulate that overfitting and publication bias give rise to the negative association between reported accuracy and sample size. Overfitting occurs when a model captures not only the underlying patterns in the training data but also its noise and idiosyncrasies, leading to poor generalization on new, unseen data. This issue is especially pronounced in situations with limited data. It can become challenging to identify overfitting in the presence of data leakage, an unintended connection between training and test datasets that arise due to mistakes in data collection and pre-processing [14, 7]. Data leakage has gained considerable attention recently in ML-based science [7], with the recognition that it likely results in overestimation of the model’s performance and unrealistic reported accuracy, particularly at small sample sizes.
A less-studied cause of overoptimism in ML-based papers is publication bias. It is known that training an ML model with limited data typically results in high variance accuracy estimates [15]. This increased variability can lead to instances where the accuracy of the model is both underestimated and overestimated. However, models with higher estimated accuracy are more likely to be published [16], a phenomenon known as publication bias or the file drawer effect in the social and medical sciences [17]. As a result, models with inflated estimates of the performance are more likely to be published. Meanwhile, models trained with larger sample sizes produce results that are more accurate and exhibit less variability, making them a more reliable measure of accuracy. Therefore, both data leakage and publication bias have a greater impact on models when the sample size is small, and this effect diminishes as the sample size increases. Hence, we posit that the negative association observed in Fig. 1 is attributable to data leakage and publication bias. Both causes of overoptimism are likely compounded by the incentives created by the academic community’s outsized emphasis on high accuracy as a primary reason for publication of new methodologies in ML-based science and the increased analytical flexibility of ML methods [18].
@@ 1.1 Proposed model
We introduce the main idea of our modeling approach here. A more detailed explanation appears in the Materials and Methods section. The relationship between training dataset size $n$ and a machine learning model’s performance $y(n)$ (say classification accuracy) is known to be effectively modeled using equations of the form $y(n)=A+\alpha n^{\beta}$ [24, 25, 26, 27, 28]. Such models, known as power laws, have found application in a variety of areas such as network science [29], social science [30] and infection disease surveillance [31].
@@
  • $\mathbf{w}_{n}\sim\mathcal{N}\left(\mu_{n},\sigma_{n}^{2}\right)$ is a random variable with a Gaussian distribution. The idea behind adding $\mathbf{w}_{n}$ is twofold. The random variable $\mathbf{w}_{n}$ accounts for the inherent variations in the estimates of the model accuracy and the effects of data leakage on estimated accuracy. Variations in model accuracy arise due to differences in data selection and data splitting for model training and testing, and they scale with $1/\sqrt{n}$ [32, 33]. We incorporate this information and use $\sigma_{n}=c_{1}/\sqrt{n}$, where $c_{1}$ is a constant, to quantitatively model the variations in model accuracy. Similarly, previous studies have shown that data leakage inflates estimated model accuracies with $n^{-0.5}$. As a result, we adjust the mean of the Gaussian distribution to $\mu_{n}=\zeta n^{-0.5}$ [34, 32] to model this. Here, $\zeta$ denotes the overfitting index, or the bias in the average accuracy estimate due to data leakage.
@@ 2 Results
Experiment 2 is a realistic simulation of the real-world process of ML model development, integrating elements of overfitting and selective reporting based on a minimal accuracy threshold. We consider binary classification problems and first empirically derive the true learning curve by progressively increasing sample sizes in ML model training. Next we model data leakage by performing feature selection on all data, then split into a training and test set for model training and evaluation. Finally, we only “report” accuracies greater than a pre-established threshold (unknown to the recovery algorithm). Our model uses the reported accuracies to estimate the learning curve. This experiment is pivotal as it showcases the model’s robustness and validity in replicating real-world ML development scenarios, where overfitting and publication bias are prevalent.
@@
Fig 2(a) and 2(b) show the results, for Problem 1 and Problem 2, respectively. The blue points represent reported accuracies that are inflated due to data leakage and publication bias. The plots further show that the estimated learning curves (red-dashdot lines) and the true learning curves (green-dashed lines) are in close agreement.
Experiment 2 (Simulating data leakage and publication bias in binary classification)
We consider two binary classification problems with high dimensional features $\mathbf{x}$, and corresponding labels $\mathbf{z}$. The detailed description of the feature vectors and their corresponding labels are provided in the SI. We consider the scenario where $K=20$ research teams train ML models, with each team performing feature selection on the complete dataset (prior to train-test split) to simulate a common form of data leakage and overfitting [12]. Then, each team randomly selected a training set (70% of the data) and trained a different ML model. Finally, each team produced an estimate of the accuracy of their classifier on the remaining (30%) of the data. To account for publication bias, only those accuracies that are greater than a pre-specified threshold $\gamma_{n}$ are reported. To establish a baseline for comparing the learning curve estimated from only the overoptimistic samples, we attain estimates of the true learning curve by iteratively increasing the sample size and training the model without data leakage and without publication bias. We repeat this process 100 times and average over these values.
@@
Figure 3: Evaluation of the proposed method per Experiment 2 where we simulate data leakage and publication bias in binary classification, (a) Problem 1, (b) Problem 2. The results show the overoptimistic accuracies (blue circles), fit to the overoptimistic results (blue line), the new estimates of the learning curve (red line) along the true learning curve (green line). The y-axis is in linear and the x-axis is in log scale.
@@ 3.1 Observation model for reported classification accuracy
Overfitting: The literature on machine learning (ML) in scientific research highlights instances of data leakage, a critical issue that leads to models overfitting the test data and producing overly optimistic estimates of classification accuracy. Data leakage often stems from improper separation between training and test datasets. A detailed classification of data leakage within ML research is presented in [7]. Furthermore, smaller datasets exacerbate the problem, as models are more likely to learn from noise and unique dataset characteristics, thus increasing the risk of overfitting. Evidence suggests that reusing a test set multiple times during model development introduces significant overfitting errors, quantifiable as $\Omega\left(tn^{-0.5}\right)$, with $t$ representing the test set’s number of reuses [34, 32]. Our proposed model, outlined in Eq. (1), integrates this understanding by incorporating Gaussian noise with an elevated mean of $\zeta n^{-0.5}$ to account for overfitting effects.
@@ Appendix C Extended results
where $\mathbf{w}_{n}\sim\mathcal{N}\left(\zeta n^{-0.5},\sigma_{n}^{2}\right)$, and $\sigma_{n}=c_{1}n^{-0.5}$, and subsequently subjected to truncation due to the selection mechanism. We categorize each ML problem as ‘easy’ if the performance limit $A>0.75$, indicating greater separability, and as ‘difficult’ if $A\leq 0.75$. Additionally, the parameter $\beta$ controls the convergence rate, with larger values in absolute value indicating faster convergence. Furthermore, higher values of the overfitting parameter $\zeta$ result in greater data leakage and overfitting to the test set. To cover a broad range of scenarios, we consider the following 7 cases:
@@
Fig. 6 shows the published classification accuracies (blue circles), and the fit to the overoptimistic results (blue) along with the estimated, debiased learning curves (red-dashdot lines) across these five fields. The dimmed blue circles in these figures are considered as outliers which have been excluded from our analysis. To identify and remove these outliers, we utilized a quantile regression approach, and discard samples below the $0.1$ quantiles. Fig. 7 shows the published results and the $0.1$ quantile. The faded blue circles indicates samples that are below this line and are considered as outliers.

### v3

@@ Abstract
Machine Learning (ML) is increasingly used across many disciplines with impressive reported results. However, recent studies suggest published performance of ML models are often overoptimistic. Validity concerns are underscored by findings of an inverse relationship between sample size and reported accuracy in published ML models, contrasting with the theory of learning curves where accuracy should improve or remain stable with increasing sample size. This paper investigates factors contributing to overoptimism in ML-driven science, focusing on overfitting and publication bias. We introduce a novel stochastic model for observed accuracy, integrating parametric learning curves and the aforementioned biases. We construct an estimator that corrects for these biases in observed data. Theoretical and empirical results show that our framework can estimate the underlying learning curve, providing realistic performance assessments from published results. Applying the model to meta-analyses of classifications of neurological conditions, we estimate the inherent limits of ML-based prediction in each domain.
@@ 1 Introduction
The authors in [8] postulate that overfitting and publication bias give rise to the negative association between reported accuracy and sample size. Overfitting occurs when a model captures not only the underlying patterns in the training data but also its noise and idiosyncrasies, leading to poor generalization on new, unseen data. This issue is especially pronounced in situations with limited data. It can become challenging to identify overfitting to the test set. This can happen due to unintended interdependencies between training and test datasets that can arise during model development [14, 7]. Model development is inherently an iterative, adaptive process [15] where researchers reuse tests sets to refine models. However, repeated use of the same test set can lead to inadvertently learning specific patterns or noise unique to that set, exacerbating overfitting issues. This problem has gained considerable attention recently in ML-based science [7], with the recognition that it likely results in overestimation of model performance and unrealistic reported accuracy, particularly at small sample sizes [16].
A less-studied cause of overoptimism in ML-based papers is publication bias. It is known that training an ML model with limited data typically results in high variance accuracy estimates [17]. This increased variability can lead to instances where the accuracy of the model is both underestimated and overestimated. However, models with higher estimated accuracy are more likely to be published [18], a phenomenon known as publication bias or the file drawer effect in the social and medical sciences [19]. As a result, models with inflated estimates of the performance are more likely to be published. Meanwhile, models trained with larger sample sizes produce results that are more accurate and exhibit less variability, making them a more reliable measure of accuracy. Therefore, both overfitting and publication bias have a greater impact on models when the sample size is small, and this effect diminishes as the sample size increases. Hence, we posit that the negative association observed in Fig. 1 is attributable to overfitting and publication bias. Both causes of overoptimism are likely compounded by the incentives created by the academic community’s outsized emphasis on high accuracy as a primary reason for publication of new methodologies in ML-based science and the increased analytical flexibility of ML methods [20].
@@ 1.1 Proposed model
We introduce the main idea of our modeling approach here. A more detailed explanation appears in the Methods section. The relationship between training dataset size $n$ and a machine learning model’s performance $y(n)$ (say classification accuracy) is known to be effectively modeled using equations of the form $y(n)=A+\alpha n^{\beta}$ [26, 27, 28, 29, 30]. Such models, known as power laws, have found application in a variety of areas such as network science [31], social science [32] and infection disease surveillance [33].
@@
  • $\mathbf{w}_{n}\sim\mathcal{N}\left(\mu_{n},\sigma_{n}^{2}\right)$ is a random variable with a Gaussian distribution. The idea behind adding $\mathbf{w}_{n}$ is twofold. The random variable $\mathbf{w}_{n}$ accounts for the inherent variations in the estimates of the model accuracy and the effects of overfitting on estimated accuracy. Variations in model accuracy arise due to differences in data selection and data splitting for model training and testing, and they scale with $1/\sqrt{n}$ [34, 35]. We incorporate this information and use $\sigma_{n}=c_{1}/\sqrt{n}$, where $c_{1}$ is a constant, to quantitatively model the variations in model accuracy. Similarly, previous studies have shown that overfitting to test set inflates estimated model accuracies with $n^{-0.5}$. As a result, we adjust the mean of the Gaussian distribution to $\mu_{n}=\zeta n^{-0.5}$ [15, 34] to model this. Here, $\zeta$ denotes the overfitting index, or the bias in the average accuracy estimate due to overfitting.
@@ 2 Results
Experiment 2 is a realistic simulation of the real-world process of ML model development, integrating elements of overfitting and selective reporting based on a minimal accuracy threshold. We consider binary classification problems and first empirically derive the true learning curve by progressively increasing sample sizes in ML model training. Next we model overfitting to the test set by performing feature selection on all data, then split into a training and test set for model training and evaluation. Finally, we only “report” accuracies greater than a pre-established threshold (unknown to the recovery algorithm). Our model uses the reported accuracies to estimate the learning curve. This experiment is pivotal as it showcases the model’s robustness and validity in replicating real-world ML development scenarios, where overfitting and publication bias are prevalent.
@@
Fig 2(a) and 2(b) show the results, for Problem 1 and Problem 2, respectively. The blue points represent reported accuracies that are inflated due to overfitting and publication bias. The plots further show that the estimated learning curves (red-dashdot lines) and the true learning curves (green-dashed lines) are in close agreement.
Experiment 2 (Simulating overfitting to the test set and publication bias in binary classification)
We consider two binary classification problems with high dimensional features $\mathbf{x}$, and corresponding labels $\mathbf{z}$. The detailed description of the feature vectors and their corresponding labels are provided in the SI. We consider the scenario where $K=20$ research teams train ML models, with each team performing feature selection on the complete dataset (prior to train-test split) to simulate a common form of overfitting [12]. Then, each team randomly selected a training set (70% of the data) and trained a different ML model. Finally, each team produced an estimate of the accuracy of their classifier on the remaining (30%) of the data. To account for publication bias, only those accuracies that are greater than a pre-specified threshold $\gamma_{n}$ are reported. To establish a baseline for comparing the learning curve estimated from only the overoptimistic samples, we attain estimates of the true learning curve by iteratively increasing the sample size and training the model without overfitting to the test set and without publication bias. We repeat this process 100 times and average over these values.
@@
Figure 3: Evaluation of the proposed method per Experiment 2 where we simulate overfitting and publication bias in binary classification, (a) Problem 1, (b) Problem 2. The results show the overoptimistic accuracies (blue circles), fit to the overoptimistic results (blue line), the new estimates of the learning curve (red line) along the true learning curve (green line). The y-axis is in linear and the x-axis is in log scale.
@@ 3.1 Observation model for reported classification accuracy
Overfitting: The literature on machine learning (ML) in scientific research highlights the issue of overfitting to the test set, which can result in overly optimistic estimates of classification accuracy. One prevalent practice contributing to this issue is adaptive data analysis [15, 43], where researchers repeatedly use the same test set to evaluate various models, features, or hyperparameters in an effort to improve model performance. This approach is widespread in the machine learning community, where maximizing performance metrics are the primary goal. However, this practice becomes problematic in small sample size regimes, as models are more susceptible to learning from noise and unique dataset characteristics, thus increasing the risk of overfitting [16]. Evidence suggests that reusing a test set multiple times during model development introduces significant overfitting errors, which optimistically scales as $\sqrt{t/n}$, with $t$ representing the test set’s number of reuses [15, 34]. Our proposed model, outlined in Eq. (1), integrates this understanding by incorporating Gaussian noise with an elevated mean of $\zeta n^{-0.5}$ to account for overfitting effects.
@@ Appendix C Extended results
where $\mathbf{w}_{n}\sim\mathcal{N}\left(\zeta n^{-0.5},\sigma_{n}^{2}\right)$, and $\sigma_{n}=c_{1}n^{-0.5}$, and subsequently subjected to truncation due to the selection mechanism. We categorize each ML problem as ‘easy’ if the performance limit $A>0.75$, indicating greater separability, and as ‘difficult’ if $A\leq 0.75$. Additionally, the parameter $\beta$ controls the convergence rate, with larger values in absolute value indicating faster convergence. Furthermore, higher values of the overfitting parameter $\zeta$ result in greater overfitting to the test set. To cover a broad range of scenarios, we consider the following 7 cases:
@@
Fig. 6 shows the published classification accuracies (blue circles), and the fit to the overoptimistic results (blue) along with the estimated, debiased learning curves (red-dashdot lines) across these five fields. The dimmed blue circles in these figures are considered as outliers which have been excluded from our analysis. To identify and remove these outliers, we utilized a quantile regression approach, and discard samples below the $0.1$ quantiles. Fig. 7 shows the published results and the $0.1$ quantile.